OpenStudy (anonymous):

4b^3-36b=0 solve for b

OpenStudy (anonymous):

4b^3-36b 4b(b^2-9) 4b(b-3)(b+3) b=1/4,3,-3

OpenStudy (anonymous):

4b(b^2 - 9) = 0 4b(b-3)(b+3) = 0 b = 0, 3, -3

OpenStudy (anonymous):

haha my bad yea that is 0 not 1/4

OpenStudy (anonymous):

4b^3-36b 4b(b^2-9) 4b(b-3)(b+3) b=1/4,3,-3